In a typical fruit drying process, fresh apricots without seeds (contains 82% of humidity) are dried in a dryer. After drying process, dry apricots contain 22% humidity. What is the mass loss of apricots during drying process. In order to obtain 1 kg dry apricots, how many kg of fresh apricots arerequired.


1
Expert's answer
2021-12-19T17:46:01-0500

Let, 100 kg of wet apricot with 82% wet humidity enters the dryer. Then, the moisture present in apricot =82 kg.

Solid content in wet apricot = 18 kg.

In the outlet humidity is 22%. Dry solid present = 78%. Now, solid content in dried apricot = solid content in wet apricot = 18kg.

So, dried apricot weight = 18kg / 0.78 = 300/13 kg. Mass loss during drying = 100 - 300/13 = 1000/13 kg. Or mass loss = 1000/13 kg / 100 kg * 100% = 76.9231% of the wet apricot entered.

In order to obtain 300/13 kg of dry apricot, fresh apricot needed = 100kg. Then for 1kg of dry apricot, fresh apricot will be needed = 100 / 300/13 = 13/3 kg = 4.333 kg